OpenStudy (anonymous):

Given: 16x2 + 25y2 - 96x - 200y + 144 = 0. Write the standard form of the ellipse. 1) Find the vertices

OpenStudy (anonymous):

pair them up first \[(16x ^{2}-96x)+(25y ^{2}-200y)+144=0\]

OpenStudy (anonymous):

\[16(x ^{2}-6x)+25(y ^{2}-8y)+144=0\]

OpenStudy (anonymous):

\[16(x ^{2}-6x+9-9)+25(y ^{2}-8y+16-16)+144=0\]

OpenStudy (anonymous):

\[16(x-3)^{2}-144+25(y-4)^{2}-400+144=0\]

OpenStudy (anonymous):

\[16(x-3)^{2}+25(y-4)^{2}=400\]

OpenStudy (anonymous):

\[(x-3)^{2}/25+(y-4)^{2}/16=1\]

OpenStudy (anonymous):

didnt get???????

OpenStudy (anonymous):

THANK YOU!